# Special pure gradings on simple Lie algebras of types $E_{6}$, $E_{7}$, $E_{8}$

### v1

@@ 5. Weyl groups of the special pure gradings on simple Lie algebras of types $E_{6}$, $E_{7}$, or $E_{8}$
Let $\mathcal{L}$ be a simple Lie algebra of type $E_{8}$ over an algebraically closed field of characteristic 00. Let $\Gamma:\mathcal{L}=\bigoplus_{g\in G}\mathcal{L}_{g}$ be the unique, up to equivalence, special pure grading with universal group $G$ and with exactly three homogeneous components: $\mathcal{L}_{s}$, $\mathcal{L}_{s^{\prime}}$ and $\mathcal{L}_{s^{\prime\prime}}$, being Cartan subalgebras. Let $W(\Gamma)$ be its Weyl group, and let $W^{\prime}(\Gamma)$ be the subgroup $W^{\prime}(\Gamma)=\{w\in W(\Gamma)\mid w(s)=s,\,w(s^{\prime})=s^{\prime},\,w(s^{\prime\prime})=s^{\prime\prime}\}$. Then the short exact sequence
@@
Let $\mathcal{L}$ be a simple Lie algebra of type $E_{7}$ over an algebraically closed field of characteristic 00. Let $\Gamma:\mathcal{L}=\bigoplus_{g\in G}\mathcal{L}_{g}$ be the unique, up to equivalence, special pure grading with universal group $G$ and with exactly three homogeneous components: $\mathcal{L}_{s}$, $\mathcal{L}_{s^{\prime}}$ and $\mathcal{L}_{s^{\prime\prime}}$, being Cartan subalgebras. Let $W(\Gamma)$ be its Weyl group, and let $W^{\prime}(\Gamma)$ be the pointwise stabilizer in $W(\Gamma)$ of $\{s,s^{\prime},s^{\prime\prime}\}$:
@@ 6. Classification of special pure gradings up to isomorphism
Recall that any special grading $\Gamma:\mathcal{L}=\bigoplus_{g\in G}\mathcal{L}_{g}$ on a semisimple Lie algebra over an algebraically closed field of characteristic 00 is almost fine ([EK25, Proposition 3.8]). If $U$ is the universal group of $\Gamma$, and we denote by $\Delta$ the grading $\Gamma$ when considered as a grading over $U$, then there is a group homomorphism $\alpha\colon U\rightarrow G$ such that $\Gamma$ is the coarsening ${}^{\alpha}\!\Delta$. This homomorphism $\alpha$ is bijective on the supports.
@@
Let $\Gamma:\mathcal{L}=\bigoplus_{g\in G}\mathcal{L}_{g}$ be a special pure grading on the simple Lie algebra $\mathcal{L}$, of type $E_{6}$, $E_{7}$, or $E_{8}$, over an algebraically closed field of characteristic 00, but now we do not assume that $G$ is the universal group of the grading. Pick an element $s$ in the support of $\Gamma$ such that $\mathcal{L}_{s}$ is a Cartan subalgebra. Denote by $T$ the subgroup of $G$ generated by the support, which is a $2$-group, as it is a quotient of the universal group, and pick a character $\chi$ of $T$ with $\chi(s)=-1$. Then $T=\ker\chi\times\langle s\rangle$. Write $\overline{T}=T/\langle s\rangle$, which is isomorphic to $\ker\chi$.
@@
Let $\Gamma:\mathcal{L}=\bigoplus_{g\in G}\mathcal{L}_{g}$ be a special pure grading on the simple Lie algebra $\mathcal{L}$ of type $E_{6}$, $E_{7}$, or $E_{8}$, over an algebraically closed field of characteristic 00. Define the sequence $\operatorname{\mathrm{Inv}}(\Gamma)$ as follows:
@@
    with $T=\langle\operatorname{\mathrm{Supp}}\Gamma\rangle$, $S=\langle t\in T\mid\mathcal{L}_{t}\ \text{is a Cartan subalgebra of $\mathcal{L}$}\rangle$ $(\simeq\mathbb{Z}_{2})$, $\overline{T}=T/S$, $H$ the subgroup of $T$ such that $\overline{H}=H/S=\overline{\pi}(\overline{E}^{\perp})$, where $E$ is the complementary lattice, and $q_{H,S}\colon\overline{T}\rightarrow\mathbb{F}_{2}$ the regular quadratic form transferred from $q|_{\overline{E}^{\perp}}$ by means of $\overline{\pi}$: $q_{H,S}\bigl{(}\overline{\pi}(x)\bigr{)}=q(x)$, for any $x\in\overline{E}^{\perp}$.
@@
Let $\Gamma:\mathcal{L}=\bigoplus_{g\in G}\mathcal{L}_{g}$ and $\Gamma^{\prime}:\mathcal{L}=\bigoplus_{g\in G}\mathcal{L}^{\prime}_{g}$ be two special pure gradings on the simple Lie algebra $\mathcal{L}$ of type $E_{6}$, $E_{7}$, or $E_{8}$, over an algebraically closed field of characteristic 00. Then $\Gamma$ is isomorphic to $\Gamma^{\prime}$ if and only if $\operatorname{\mathrm{Inv}}(\Gamma)=\operatorname{\mathrm{Inv}}(\Gamma^{\prime})$.
@@
For the converse, let us start with a simple case. Assume that $\Gamma$ and $\Gamma^{\prime}$ are equivalent to $\Gamma_{E_{8}}^{6}$ and $\operatorname{\mathrm{Inv}}(\Gamma)=\operatorname{\mathrm{Inv}}(\Gamma^{\prime})=(T,K,S)$. As in Theorem 5.7, the universal group of $\Gamma_{E_{8}}^{6}$ is $U=H\times\mathbb{F}_{2}^{2}$, and let $\alpha,\beta\colon U\rightarrow G$ be the corresponding group homomorphisms so that $\Gamma={}^{\alpha}(\Gamma_{E_{8}}^{6})$ and $\Gamma^{\prime}={}^{\beta}(\Gamma_{E_{8}}^{6})$. Then $\alpha$ and $\beta$ are isomorphisms when considered as maps $U\rightarrow T=\langle\operatorname{\mathrm{Supp}}\Gamma\rangle=\langle\operatorname{\mathrm{Supp}}\Gamma^{\prime}\rangle$. Then the composition $\beta^{-1}\circ\alpha\colon U\rightarrow U$ makes sense and it is an automorphism that preserves the flag in Theorem 5.7. Therefore, $\beta^{-1}\circ\alpha$ belongs to the Weyl group $W(\Gamma_{E_{8}}^{6})$. It follows that $\alpha=\beta\circ\omega$ for an element $\omega\in W(\Gamma_{E_{8}}^{6})$, proving that $\Gamma$ and $\Gamma^{\prime}$ are isomorphic.
@@
as in (5.5), and thus $\Gamma$ and $\Gamma^{\prime}$ are isomorphic. The converse is clear.

### v2

@@ 5. Weyl groups of the special pure gradings on simple Lie algebras of types $E_{6}$, $E_{7}$, or $E_{8}$
Let $\mathcal{L}$ be a simple Lie algebra of type $E_{8}$ over an algebraically closed field of characteristic 0. Let $\Gamma:\mathcal{L}=\bigoplus_{g\in G}\mathcal{L}_{g}$ be the unique, up to equivalence, special pure grading with universal group $G$ and with exactly three homogeneous components: $\mathcal{L}_{s}$, $\mathcal{L}_{s^{\prime}}$ and $\mathcal{L}_{s^{\prime\prime}}$, being Cartan subalgebras. Let $W(\Gamma)$ be its Weyl group, and let $W^{\prime}(\Gamma)$ be the subgroup $W^{\prime}(\Gamma)=\{w\in W(\Gamma)\mid w(s)=s,\,w(s^{\prime})=s^{\prime},\,w(s^{\prime\prime})=s^{\prime\prime}\}$. Then the short exact sequence
@@
Let $\mathcal{L}$ be a simple Lie algebra of type $E_{7}$ over an algebraically closed field of characteristic 0. Let $\Gamma:\mathcal{L}=\bigoplus_{g\in G}\mathcal{L}_{g}$ be the unique, up to equivalence, special pure grading with universal group $G$ and with exactly three homogeneous components: $\mathcal{L}_{s}$, $\mathcal{L}_{s^{\prime}}$ and $\mathcal{L}_{s^{\prime\prime}}$, being Cartan subalgebras. Let $W(\Gamma)$ be its Weyl group, and let $W^{\prime}(\Gamma)$ be the pointwise stabilizer in $W(\Gamma)$ of $\{s,s^{\prime},s^{\prime\prime}\}$:
@@ 6. Classification of special pure gradings up to isomorphism
Recall that any special grading $\Gamma:\mathcal{L}=\bigoplus_{g\in G}\mathcal{L}_{g}$ on a semisimple Lie algebra over an algebraically closed field of characteristic 0 is almost fine ([EK25, Proposition 3.8]). If $U$ is the universal group of $\Gamma$, and we denote by $\Delta$ the grading $\Gamma$ when considered as a grading over $U$, then there is a group homomorphism $\alpha\colon U\rightarrow G$ such that $\Gamma$ is the coarsening $^{\alpha}\!\Delta$. This homomorphism $\alpha$ is bijective on the supports.
@@
Let $\Gamma:\mathcal{L}=\bigoplus_{g\in G}\mathcal{L}_{g}$ be a special pure grading on the simple Lie algebra $\mathcal{L}$, of type $E_{6}$, $E_{7}$, or $E_{8}$, over an algebraically closed field of characteristic 0, but now we do not assume that $G$ is the universal group of the grading. Pick an element $s$ in the support of $\Gamma$ such that $\mathcal{L}_{s}$ is a Cartan subalgebra. Denote by $T$ the subgroup of $G$ generated by the support, which is a $2$-group, as it is a quotient of the universal group, and pick a character $\chi$ of $T$ with $\chi(s)=-1$. Then $T=\ker\chi\times\langle s\rangle$. Write $\overline{T}=T/\langle s\rangle$, which is isomorphic to $\ker\chi$.
@@
Let $\Gamma:\mathcal{L}=\bigoplus_{g\in G}\mathcal{L}_{g}$ be a special pure grading on the simple Lie algebra $\mathcal{L}$ of type $E_{6}$, $E_{7}$, or $E_{8}$, over an algebraically closed field of characteristic 0. Define the sequence $\operatorname{\mathrm{Inv}}(\Gamma)$ as follows:
@@
    with $T=\langle\operatorname{\mathrm{Supp}}\Gamma\rangle$, $S=\langle t\in T\mid\mathcal{L}_{t}\ \text{is a Cartan subalgebra of $\mathcal{L}$}\rangle$ $(\simeq\mathbb{Z}_{2})$, $\overline{T}=T/S$, $H$ the subgroup of $T$ such that $\overline{H}=H/S=\overline{\pi}(\overline{E}^{\perp})$, where $E$ is the complementary lattice, and $q_{H,S}\colon\overline{H}\rightarrow\mathbb{F}_{2}$ the regular quadratic form transferred from $q|_{\overline{E}^{\perp}}$ by means of $\overline{\pi}$: $q_{H,S}\bigl(\overline{\pi}(x)\bigr)=q(x)$, for any $x\in\overline{E}^{\perp}$.
@@
Let $\Gamma:\mathcal{L}=\bigoplus_{g\in G}\mathcal{L}_{g}$ and $\Gamma^{\prime}:\mathcal{L}=\bigoplus_{g\in G}\mathcal{L}^{\prime}_{g}$ be two special pure gradings on the simple Lie algebra $\mathcal{L}$ of type $E_{6}$, $E_{7}$, or $E_{8}$, over an algebraically closed field of characteristic 0. Then $\Gamma$ is isomorphic to $\Gamma^{\prime}$ if and only if $\operatorname{\mathrm{Inv}}(\Gamma)=\operatorname{\mathrm{Inv}}(\Gamma^{\prime})$.
@@
[…]
@@
as in (5.5), and thus $\Gamma$ and $\Gamma^{\prime}$ are isomorphic.